### v2

@@ 7 Separable utilities are $\Sigma$-Gale-substitutes
In this section, we prove Theorem 1.4. See 1.4 Recall the following simple monotonicity property of supergradients of univariate concave functions.
Lemma 7.1.
@@
Let $v\,:\,\mathbb{R}\to\mathbb{R}$ be a univariate concave function, $\alpha,\alpha^{\prime}\in\operatorname{dom}(f)$, and $g\in\partial^{\wedge}v(\alpha)$, $g^{\prime}\in\partial^{\wedge}v(\alpha^{\prime})$. Then, $\alpha<\alpha^{\prime}$ implies $g\geq g^{\prime}$. Conversely, if $g>g^{\prime}$, then $\alpha\leq\alpha^{\prime}$.
We now adapt Lemma 5.3 to separable utilities.
Lemma 7.2.
Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function. For any $b>0$, $\mathbf{q}\in\operatorname{pdom}(u)$, $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, we have $u(y)>0$. Further, $y\in\mathcal{GD}^{u}(b,\mathbf{q})$ if and only if there exist supergradients $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$, $j\in{\mathcal{M}}$ satisfying
  1. (i)
    $g_{j}\geq 0$; and
  2. (ii)
    $g_{j}\leq q_{j}u(y)/b$ with equality whenever $y_{j}>0$.
@@
The proof of Theorem 1.4 follows by the next three lemmas, that prove the three properties in Definition 2.14 for separable utilities: property (i) in Lemma 7.3, property (ii) in Lemma 7.4, and property ($\Sigma$) in Lemma 7.5. A similar argument was used in Lemma 7.1 in Garg et al. (2022a) to show a variant of the Gale substitutes property for capped SPLC utilities.
Lemma 7.3.
@@
Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, and let $b>0$, $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ and $y\in\mathcal{GD}^{u}(\mathbf{q},b)$. If $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$.
@@
Let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\prime}_{j}=q_{j}\}$. Let us select $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ minimizing the potential $\Psi(y^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in J}(y^{\prime}_{j}-y_{j})^{+}$. Our goal is to show $\Psi(y^{\prime})=0$. Throughout the proof, let $g^{\prime}_{j}\in\partial^{\wedge}v_{j}(y^{\prime}_{j})$ and $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$ as in Lemma 7.2.
First, we show $u(y^{\prime})\geq u(y)$. For a contradiction, assume $u(y^{\prime})<u(y)$. Then, according to Lemma 7.2, for every $j\in{\mathcal{M}}$ with $y_{j}>0$ and $q_{j}>0$ we have
$g^{\prime}_{j}\leq q^{\prime}_{j}u(y^{\prime})/b<q_{j}u(y)/b=g_{j}\,,$
while for $j\in{\mathcal{M}}$ with $q_{j}=0$ we have $g^{\prime}_{j}=g_{j}=0$. Consequently, for each $y_{j}>0$, if $q_{j}>0$ then Lemma 7.1 implies $y^{\prime}_{j}\geq y_{j}$, and if $q_{j}=0$ then $v_{j}(y_{j})=v_{j}(y_{j}^{\prime})$, both being equal to the maximum value of $v_{j}(.)$. These imply $u(y^{\prime})\geq u(y)$, a contradiction.
@@
$g^{\prime}_{j}=\frac{q^{\prime}_{j}}{b}u(y^{\prime})\geq\frac{q_{j}}{b}u(y)\geq g_{j}\,.$ (7)
@@
Let us first consider the case $u(y^{\prime})>u(y)$. For all $j\in J$ with $q^{\prime}_{j}=q_{j}>0$, we get $g^{\prime}_{j}>g_{j}$. By Lemma 7.1, $y^{\prime}_{j}\leq y_{j}$ follows for such goods $j$. On the other hand, if $q^{\prime}_{j}=q_{j}=0$, then this implies $g^{\prime}_{j}=g_{j}=0$. For such indices, if $y^{\prime}_{j}>y_{j}$, then replacing $y^{\prime}_{j}$ by $y_{j}$ results in another bundle $y^{\prime\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ with $\Psi(y^{\prime\prime})<\Psi(y^{\prime})$, a contradiction.
@@
For the rest of the proof, let us assume that $u(y^{\prime})=u(y)$. We next show $y^{\prime}_{j}\geq y_{j}$ for all $j\in{\mathcal{M}}\setminus J$, that is, for all goods with $q^{\prime}_{j}<q_{j}$. This is clearly true if $y_{j}=0$. Whenever $y_{j}>0$ for such an index, we have
$g^{\prime}_{j}\leq\frac{q^{\prime}_{j}}{b}u(y^{\prime})<\frac{q_{j}}{b}u(y)=g_{j}\,,$
@@
and therefore $y^{\prime}_{j}\geq y_{j}$ by Lemma 7.1. By monotonicity, $v_{j}(y^{\prime}_{j})\geq v_{j}(y_{j})$ for all $j\in{\mathcal{M}}\setminus J$. Since $u(y^{\prime})=u(y)$, we have $\sum_{j\in J}v_{j}(y^{\prime}_{j})\leq\sum_{j\in J}v_{j}(y_{j})$.
@@
From (7), it follows that $g^{\prime}_{j}=g_{j}$ for all $j\in J$ with $y^{\prime}_{j}>0$. This is because the first inequality is now equality, and the second inequality may only be strict if $y_{j}=0<y^{\prime}_{j}$. By Lemma 7.1, $g_{j}\geq g^{\prime}_{j}$ in this case, leading to a contradiction.
@@
Assume for a contradiction $\Psi(y^{\prime})>0$, and let us select an index $k\in J$ with $y^{\prime}_{k}>y_{k}$. If $v_{k}(y^{\prime}_{k})=v_{k}(y_{k})$, then we get a contradiction, because changing $y^{\prime}_{k}$ to $y_{k}$ would preserve the the optimality conditions in Lemma 7.2, while decreasing the potential $\Psi(y^{\prime})$. Thus, $v_{k}(y^{\prime}_{k})>v_{k}(y_{k})$. By $\sum_{j\in J}v_{j}(y^{\prime}_{j})\leq\sum_{j\in J}v_{j}(y_{j})$, there must be an index $j\in J$ with $v_{j}(y^{\prime}_{j})<v_{j}(y_{j})$. We can now find suitable $\varepsilon_{k},\varepsilon_{j}>0$ such that $v_{k}(y^{\prime}_{k}-\varepsilon_{k})+v_{j}(y^{\prime}_{j}+\varepsilon_{j})=v_{k}(y^{\prime}_{k})+v_{j}(y^{\prime}_{j})$, and $y^{\prime}_{j}+\varepsilon_{j}\leq y_{j}$. This modified solution $y^{\prime\prime}$ satisfies the optimality conditions but $\Psi(y^{\prime\prime})<\Psi(y^{\prime})$, again leading to a contradiction. ∎
Lemma 7.4.
@@
Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$. Then, for any $y\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $y^{\prime}\leq y$.
@@
We argue similarly as in the previous proof: we pick $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ which minimizes $\Psi(y^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in J}(y^{\prime}_{j}-y_{j})^{+}$, and show that $\Psi(y^{\prime})=0$.
First, we show that ${u(y)}/{b}\leq{u(y^{\prime})}/{b^{\prime}}$. For a contradiction, suppose ${u(y)}/{b}>{u(y^{\prime})}/{b^{\prime}}$. If $y_{j}>0$, then, by Lemma 7.2
$g_{j}=\frac{q_{j}u(y)}{b}>\frac{q_{j}u(y^{\prime})}{b^{\prime}}\geq g^{\prime}_{j},$
which implies $v_{j}(y_{j})\leq v_{j}(y^{\prime}_{j})$ whenever $y_{j}>0$ (Lemma 7.1), and therefore, $u(y)\leq u(y^{\prime})$. Together with $b>b^{\prime}$, this implies ${u(y)}/{b}\leq{u(y^{\prime})}/{b^{\prime}}$, a contradiction.
@@
$g^{\prime}_{j}=\frac{q_{j}u(y^{\prime})}{b^{\prime}}\geq\frac{q_{j}u(y)}{b}\geq g_{j}\,.$ (8)
Assume first $u(y^{\prime})/b^{\prime}>u(y)/b$. If $q_{j}>0$, then this inequality is strict, and therefore $y^{\prime}_{j}\leq y_{j}$ for all $j\in{\mathcal{M}}$ follows by Lemma 7.1. If $q_{j}=0$, then $g^{\prime}_{j}=g_{j}=0$. If $y^{\prime}_{j}>y_{j}$ for such a good $j$, we can decrease $y^{\prime}_{j}$ to $y_{j}$ and thereby get an optimal allocation with strictly smaller $\Psi(y^{\prime})$ value, leading to a contradiction. This completes the proof for $u(y^{\prime})/b^{\prime}>u(y)/b$.
For the rest, we may assume $u(y^{\prime})/b^{\prime}=u(y)/b$. Now, $g^{\prime}_{j}=g_{j}$ must hold whenever $y^{\prime}_{j}>0$. This is because the first inequality in (8) is equality, and if the second is strict, then $y_{j}=0<y^{\prime}_{j}$ would give a contradiction to $g^{\prime}_{j}>g_{j}$.
Assume for a contradiction $\Psi(y^{\prime})>0$, and let us pick $k\in{\mathcal{M}}$ with $y^{\prime}_{k}>y_{k}$. The equality $u(y^{\prime})/b^{\prime}=u(y)/b$ and $b^{\prime}<b$ gives $u(y^{\prime})<u(y)$. We can derive a contradiction by decreasing $y^{\prime}_{k}$ and potentially increasing an $y^{\prime}_{j}$ similarly as in the proof of Lemma 7.3. ∎
Lemma 7.5.
Let $u(x)=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function such that $\mathbf{0}\in\operatorname{pdom}(u)$. Then, for any $\mathbf{q}\geq\mathbf{0}$, $b>0$ and any $y\in\mathcal{GD}^{u}(\mathbf{0},b)$, there exists a $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $y^{\prime}\leq y$.
@@
$\mathbf{0}\in\operatorname{pdom}(u)$ implies $\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(y)<\infty$. Therefore, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$; in particular, $\mathbf{q}\in\operatorname{pdom}(u)$. Let us pick $y^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$ is minimal.
Note that $u(y)=\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(y)$. In particular, for any $\alpha\geq y_{j}$, $v_{j}(\alpha)=v_{j}(y_{j})$. For a contradiction, assume there exists a $k\in{\mathcal{M}}$ with $y_{k}<y^{\prime}_{k}$. Then, reducing $y^{\prime}_{k}$ to $y_{k}$ may only increase $b\log u(y^{\prime})-\left\langle\mathbf{q},y^{\prime}\right\rangle$, while decreasing $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$. This contradicts the choice of $y^{\prime}$. ∎
8 Generalized network utilities
In this section, we prove Theorem 1.5. We will first need a characterization of subgradients by linear programming duality. We state this in a more general form. The proof based on LP duality is deferred to Appendix B.
Lemma 8.1.
Let $\mathbf{A}\in\mathbb{R}^{k\times\ell}$, $\mathbf{B}\in\mathbb{R}^{k\times m}$, $c\in\mathbb{R}^{\ell}$, $b\in\mathbb{R}^{k}$. Let $\mathcal{D}\subseteq\mathbb{R}^{m}$ be the set of vectors $x\in\mathbb{R}^{m}$ such that the following parametric linear program is feasible and bounded, and let us define $\varphi:\mathcal{D}\to\mathbb{R}$ such that $\varphi(x)$ is the optimum value of this program:
$\max\leavevmode\nobreak$ $c^{\top}z$ ($\operatorname{Primal}(x)$)
$\mathbf{A}z$ $\leq\mathbf{B}x+b$
$z$ $\geq 0$
Then, for every $x\in\mathcal{D}$, $g\in\partial^{\wedge}\varphi(x)$ if and only if $g=\mathbf{B}^{\top}\pi$ for an optimal solution $\pi\in\mathbb{R}_{\geq 0}^{k}$ to the dual linear program
$\min\leavevmode\nobreak$ $x^{\top}\mathbf{B}^{\top}\pi+b^{\top}\pi$ ($\operatorname{Dual}(x)$)
$\mathbf{A}^{\top}\pi$ $\geq c$
$\pi$ $\geq 0\,.$
Consider a generalized flow instance defined in a graph $G=(V,E)$, sink $t\in V$, capacities $h\in\mathbb{R}_{\geq 0}^{E}$, gain factors $\gamma\in\mathbb{R}_{>0}^{E}$ and supply set $S={\mathcal{M}}$. For the rest of this section, let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function defined in the form $\operatorname{GenflowLP}(x)$ in Section 2.
Using Lemma 8.1 and complementary slackness, we get the following characterization of subgradients of generalized network utility functions. Note that for a generalized network utility function, the domain is $\mathcal{D}=\mathbb{R}_{\geq 0}^{\mathcal{M}}$, because $f=0$ is always a feasible solution to $\operatorname{GenflowLP}(x)$, and the definition assumes that the objective is always bounded.
Lemma 8.2.

### v3

@@ Appendix C Separable utilities are $\Sigma$-Gale-substitutes
In this Appendix, we prove Theorem 1.4 for the general case. Recall the following simple monotonicity property of supergradients of univariate concave functions.
Lemma C.1.
@@
Let $v\,:\,\mathbb{R}\to\mathbb{R}$ be a univariate concave function, $\alpha,\alpha^{\prime}\in\operatorname{dom}(v)$, and $g\in\partial^{\wedge}v(\alpha)$, $g^{\prime}\in\partial^{\wedge}v(\alpha^{\prime})$. Then, $\alpha<\alpha^{\prime}$ implies $g\geq g^{\prime}$. Conversely, if $g>g^{\prime}$, then $\alpha\leq\alpha^{\prime}$.
@@
The proof of Theorem 1.4 follows by the next three lemmas, that prove the three properties in Definition 2.14 for separable utilities: property (i) in Lemma C.2, property (ii) in Lemma C.3, and property ($\Sigma$) in Lemma C.4. A similar argument was used in Lemma 7.1 in Garg et al. (2022a) to show a variant of the Gale substitutes property for capped SPLC utilities.
Lemma C.2.
@@
Let $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, and let $b>0$, $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ and $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$. If $\mathbf{q}\geq\mathbf{q}^{\prime}$, then there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $y^{\prime}_{j}\leq y_{j}$ whenever $q_{j}^{\prime}=q_{j}$.
@@
Let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\prime}_{j}=q_{j}\}$. Let us select $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ minimizing the potential $\Psi(\mathbf{y}^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in J}(y^{\prime}_{j}-y_{j})^{+}$. Our goal is to show $\Psi(\mathbf{y}^{\prime})=0$. Throughout the proof, let $g^{\prime}_{j}\in\partial^{\wedge}v_{j}(y^{\prime}_{j})$ and $g_{j}\in\partial^{\wedge}v_{j}(y_{j})$ as in Lemma 6.1.
First, we show $u(\mathbf{y}^{\prime})\geq u(\mathbf{y})$. For a contradiction, assume $u(\mathbf{y}^{\prime})<u(\mathbf{y})$. Then, according to Lemma 6.1, for every $j\in{\mathcal{M}}$ with $y_{j}>0$ and $q_{j}>0$ we have
$g^{\prime}_{j}\leq q^{\prime}_{j}u(\mathbf{y}^{\prime})/b<q_{j}u(\mathbf{y})/b=g_{j}\,$,
while for $j\in{\mathcal{M}}$ with $q_{j}=0$ we have $g^{\prime}_{j}=g_{j}=0$. Consequently, for each $y_{j}>0$, if $q_{j}>0$ then Lemma C.1 implies $y^{\prime}_{j}\geq y_{j}$, and if $q_{j}=0$ then $v_{j}(y_{j})=v_{j}(y_{j}^{\prime})$, both being equal to the maximum value of $v_{j}(.)$. These imply $u(\mathbf{y}^{\prime})\geq u(\mathbf{y})$, a contradiction.
@@
$g^{\prime}_{j}=\frac{q^{\prime}_{j}}{b}u(\mathbf{y}^{\prime})\geq\frac{q_{j}}{b}u(\mathbf{y})\geq g_{j}\,.$ (7)
@@
Let us first consider the case $u(\mathbf{y}^{\prime})>u(\mathbf{y})$. For all $j\in J$ with $q^{\prime}_{j}=q_{j}>0$, we get $g^{\prime}_{j}>g_{j}$. By Lemma C.1, $y^{\prime}_{j}\leq y_{j}$ follows for such goods $j$. On the other hand, if $q^{\prime}_{j}=q_{j}=0$, then this implies $g^{\prime}_{j}=g_{j}=0$. For such indices, if $y^{\prime}_{j}>y_{j}$, then replacing $y^{\prime}_{j}$ by $y_{j}$ results in another bundle $\mathbf{y}^{\prime\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ with $\Psi(\mathbf{y}^{\prime\prime})<\Psi(\mathbf{y}^{\prime})$, a contradiction.
@@
For the rest of the proof, let us assume that $u(\mathbf{y}^{\prime})=u(\mathbf{y})$. We next show $y^{\prime}_{j}\geq y_{j}$ for all $j\in{\mathcal{M}}\setminus J$, that is, for all goods with $q^{\prime}_{j}<q_{j}$. This is clearly true if $y_{j}=0$. Whenever $y_{j}>0$ for such an index, we have
$g^{\prime}_{j}\leq\frac{q^{\prime}_{j}}{b}u(\mathbf{y}^{\prime})<\frac{q_{j}}{b}u(\mathbf{y})=g_{j}\,$,
@@
and therefore $y^{\prime}_{j}\geq y_{j}$ by Lemma C.1. By monotonicity, $v_{j}(y^{\prime}_{j})\geq v_{j}(y_{j})$ for all $j\in{\mathcal{M}}\setminus J$. Since $u(\mathbf{y}^{\prime})=u(\mathbf{y})$, we have $\sum_{j\in J}v_{j}(y^{\prime}_{j})\leq\sum_{j\in J}v_{j}(y_{j})$.
@@
From (7), it follows that $g^{\prime}_{j}=g_{j}$ for all $j\in J$ with $y^{\prime}_{j}>0$. This is because the first inequality is now equality, and the second inequality may only be strict if $y_{j}=0<y^{\prime}_{j}$. By Lemma C.1, $g_{j}\geq g^{\prime}_{j}$ in this case, leading to a contradiction. It also follows that $g^{\prime}_{k}=g_{k}$ for all $k\in J$ with $y_{k}>0$ from the same argument.
@@
Assume for a contradiction $\Psi(\mathbf{y}^{\prime})>0$, and let us select an index $j\in J$ with $y^{\prime}_{j}>y_{j}$. If $v_{j}(y^{\prime}_{j})=v_{j}(y_{j})$, then we get a contradiction, because changing $y^{\prime}_{j}$ to $y_{j}$ would preserve the optimality conditions in Lemma 6.1, while decreasing the potential $\Psi(\mathbf{y}^{\prime})$. Thus, $v_{j}(y^{\prime}_{j})>v_{j}(y_{j})$. By $\sum_{j^{\prime}\in J}v_{j^{\prime}}(y^{\prime}_{j^{\prime}})\leq\sum_{j^{\prime}\in J}v_{j^{\prime}}(y_{j^{\prime}})$, there must be an index $k\in J$ with $v_{k}(y^{\prime}_{k})<v_{k}(y_{k})$. We note as $u(\mathbf{y})=u(\mathbf{y}^{\prime})$ and $j\in J$ and $k\in J$, this implies $g^{\prime}_{k}=g_{k}=\frac{q^{\prime}_{k}}{b}u(\mathbf{y}^{\prime})$ and $g^{\prime}_{j}=g_{j}=\frac{q^{\prime}_{j}}{b}u(\mathbf{y}^{\prime})$. We can now find small $\varepsilon_{k},\varepsilon_{j}>0$ such that $\varepsilon_{k}\cdot q^{\prime}_{k}=\varepsilon_{j}\cdot q^{\prime}_{j}$; $y^{\prime}_{j}-\varepsilon_{j}\geq y_{j}$ and $y^{\prime}_{k}+\varepsilon_{k}\leq y_{k}$. We note after making the update $y^{\prime\prime}_{j}=y^{\prime}_{j}-\varepsilon_{j}\geq y_{j}$ and $y^{\prime\prime}_{k}=y^{\prime}_{k}+\varepsilon_{k}\leq y_{k}$. Additionally, $v_{k}(y^{\prime}_{k}+\varepsilon_{k})+v_{j}(y^{\prime}_{j}-\varepsilon_{j})=v_{k}(y^{\prime}_{k})+v_{j}(y^{\prime}_{j})$ and $y^{\prime\prime}_{j}q^{\prime}_{j}+y^{\prime\prime}_{k}q^{\prime}_{k}=y^{\prime}_{j}q^{\prime}_{j}+y^{\prime}_{k}q^{\prime}_{k}$. Therefore, this modified solution $\mathbf{y}^{\prime\prime}$ satisfies the optimality conditions but $\Psi(\mathbf{y}^{\prime\prime})<\Psi(\mathbf{y}^{\prime})$, again leading to a contradiction. ∎
Lemma C.3.
@@
Let $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$. Then, for any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{q},b)$, there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $\mathbf{y}^{\prime}\leq\mathbf{y}$.
@@
We argue similarly as in the previous proof: we pick $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ which minimizes $\Psi(\mathbf{y}^{\prime})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{j\in{\mathcal{M}}}(y^{\prime}_{j}-y_{j})^{+}$, and show that $\Psi(\mathbf{y}^{\prime})=0$.
First, we show that ${u(\mathbf{y})}/{b}\leq{u(\mathbf{y}^{\prime})}/{b^{\prime}}$. For a contradiction, suppose ${u(\mathbf{y})}/{b}>{u(\mathbf{y}^{\prime})}/{b^{\prime}}$. If $y_{j}>0$, then, by Lemma 6.1
$g_{j}=\frac{q_{j}u(\mathbf{y})}{b}>\frac{q_{j}u(\mathbf{y}^{\prime})}{b^{\prime}}\geq g^{\prime}_{j},$
which implies $v_{j}(y_{j})\leq v_{j}(y^{\prime}_{j})$ whenever $y_{j}>0$ (Lemma C.1), and therefore, $u(\mathbf{y})\leq u(\mathbf{y}^{\prime})$. Together with $b>b^{\prime}$, this implies ${u(\mathbf{y})}/{b}\leq{u(\mathbf{y}^{\prime})}/{b^{\prime}}$, a contradiction.
@@
$g^{\prime}_{j}=\frac{q_{j}u(\mathbf{y}^{\prime})}{b^{\prime}}\geq\frac{q_{j}u(\mathbf{y})}{b}\geq g_{j}\,.$ (8)
Assume first $u(\mathbf{y}^{\prime})/b^{\prime}>u(\mathbf{y})/b$. If $q_{j}>0$, then inequality (8) is strict, and therefore $y^{\prime}_{j}\leq y_{j}$ for all $j\in{\mathcal{M}}$ follows by Lemma C.1. If $q_{j}=0$, then $g^{\prime}_{j}=g_{j}=0$. If $y^{\prime}_{j}>y_{j}$ for such a good $j$, we can decrease $y^{\prime}_{j}$ to $y_{j}$ and thereby get an optimal allocation with strictly smaller $\Psi(\mathbf{y}^{\prime})$ value, leading to a contradiction. This completes the proof for $u(\mathbf{y}^{\prime})/b^{\prime}>u(\mathbf{y})/b$.
For the rest, we may assume $u(\mathbf{y}^{\prime})/b^{\prime}=u(\mathbf{y})/b$. Now, $g^{\prime}_{j}=g_{j}$ must hold whenever $y^{\prime}_{j}>0$. This is because the first inequality in (8) is equality, and if the second is strict, then $y_{j}=0<y^{\prime}_{j}$ would give a contradiction to $g^{\prime}_{j}>g_{j}$. It also follows from the same argument that $g^{\prime}_{k}=g_{k}$ for all $k\in{\mathcal{M}}$ with $y_{k}>0$.
Assume for a contradiction that $\Psi(\mathbf{y}^{\prime})>0$, and let us select an index $j\in{\mathcal{M}}$ such that $y^{\prime}_{j}>y_{j}$. If $v_{j}(y^{\prime}_{j})=v_{j}(y_{j})$, we obtain a contradiction because changing $y^{\prime}_{j}$ to $y_{j}$ would preserve the optimality conditions in Lemma 6.1 while decreasing the potential $\Psi(\mathbf{y}^{\prime})$. Hence, we must have $v_{j}(y^{\prime}_{j})>v_{j}(y_{j})$. Since $\sum_{j^{\prime}\in{\mathcal{M}}}v_{j^{\prime}}(y^{\prime}_{j^{\prime}})\leq\sum_{j^{\prime}\in{\mathcal{M}}}v_{j^{\prime}}(y_{j^{\prime}})$, there must exist an index $k\in{\mathcal{M}}$ with $v_{k}(y^{\prime}_{k})<v_{k}(y_{k})$. Noting that $u(\mathbf{y})=u(\mathbf{y}^{\prime})$ and $j,k\in{\mathcal{M}}$, this implies $g^{\prime}_{k}=g_{k}=\frac{q^{\prime}_{k}}{b}u(\mathbf{y}^{\prime})$ and $g^{\prime}_{j}=g_{j}=\frac{q^{\prime}_{j}}{b}u(\mathbf{y}^{\prime})$. We can now choose small $\varepsilon_{k},\varepsilon_{j}>0$ such that $\varepsilon_{k}\cdot q^{\prime}_{k}=\varepsilon_{j}\cdot q^{\prime}_{j}$; $y^{\prime}_{j}-\varepsilon_{j}\geq y_{j}$ and $y^{\prime}_{k}+\varepsilon_{k}\leq y_{k}$. After making this update, define $y^{\prime\prime}_{j}=y^{\prime}_{j}-\varepsilon_{j}\geq y_{j}$ and $y^{\prime\prime}_{k}=y^{\prime}_{k}+\varepsilon_{k}\leq y_{k}$. Moreover, we have $v_{k}(y^{\prime}_{k}+\varepsilon_{k})+v_{j}(y^{\prime}_{j}-\varepsilon_{j})=v_{k}(y^{\prime}_{k})+v_{j}(y^{\prime}_{j})$ and $y^{\prime\prime}_{j}q^{\prime}_{j}+y^{\prime\prime}_{k}q^{\prime}_{k}=y^{\prime}_{j}q^{\prime}_{j}+y^{\prime}_{k}q^{\prime}_{k}$. Thus, the modified solution $\mathbf{y}^{\prime\prime}$ satisfies the optimality conditions, but $\Psi(\mathbf{y}^{\prime\prime})<\Psi(\mathbf{y}^{\prime})$, again leading to a contradiction. ∎
Lemma C.4.
Let $u(\mathbf{x})=\sum_{j\in{\mathcal{M}}}v_{j}(x_{j})$ be a separable utility function such that $\mathbf{0}\in\operatorname{pdom}(u)$. Then, for any $\mathbf{q}\geq\mathbf{0}$, $b>0$ and any $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{0},b)$, there exists a $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $\mathbf{y}^{\prime}\leq\mathbf{y}$.
@@
$\mathbf{0}\in\operatorname{pdom}(u)$ implies $\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(\mathbf{y})<\infty$. Therefore, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$; in particular, $\mathbf{q}\in\operatorname{pdom}(u)$. Let us pick $\mathbf{y}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ such that $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$ is minimal.
Note that $u(\mathbf{y})=\sup_{\mathbf{y}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(\mathbf{y})$. In particular, for any $\alpha\geq y_{j}$, $v_{j}(\alpha)=v_{j}(y_{j})$. For a contradiction, assume there exists a $k\in{\mathcal{M}}$ with $y_{k}<y^{\prime}_{k}$. Then, reducing $y^{\prime}_{k}$ to $y_{k}$ may only increase $b\log u(\mathbf{y}^{\prime})-\left\langle\mathbf{q},\mathbf{y}^{\prime}\right\rangle$, while decreasing $\sum_{j\in{\mathcal{M}}}y^{\prime}_{j}$. This contradicts the choice of $\mathbf{y}^{\prime}$. ∎
Appendix D Generalized network utilities
In this Appendix, we prove Theorem 1.5 and Theorem 1.6. We will first need a characterization of subgradients by linear programming duality. We state this in a more general form.
Lemma D.1.
Let $A\in\mathbb{R}^{k\times\ell}$, $B\in\mathbb{R}^{k\times m}$, $\mathbf{c}\in\mathbb{R}^{\ell}$, $\mathbf{b}\in\mathbb{R}^{k}$. Let $\mathcal{D}\subseteq\mathbb{R}^{m}$ be the set of vectors $\mathbf{x}\in\mathbb{R}^{m}$ such that the following parametric linear program is feasible and bounded, and let us define $\varphi:\mathcal{D}\to\mathbb{R}$ such that $\varphi(\mathbf{x})$ is the optimum value of this program: